# Stochastic Gravitational Wave Background from Highly-Eccentric Stellar-Mass Binaries in the Milli-hertz Band

### v1

@@ II.2.2 Stochastic Background Calculation
here $\dot{\epsilon}(f(1+z))$ is the emitted energy spectrum of the GW sources (i.e., energy per unit comoving volume, unit proper time, and unit emitted frequency).
We note that, in general, there exists a straightforward relationship between the spectrum of the gravitational wave background produced by a cosmological distribution of discrete gravitational wave sources and the present-day comoving number density of remnants, as described by Eq. 5 in Ref. [104]. Realistic examples and further discussions on this relationship can be found in studies such as Refs. [105, 106]. For our cases, however, since the local population can have a non-negligible contribution to the overall GW background level, we include the inhomogeneous spatial distribution and non-equilibrium star formation history in the MCMC simulations afterward (see Section III.1). Therefore, in order to keep consistent, we adopt Equations (20) and (21) as a robust approach to estimate the GW background level across all channels, including cosmological background calculations.
@@ III.1 Configuration of the Simulations
As a proof of concept, we adopt the population model in Xuan et al. [33] to generate the parameters of highly eccentric, stellar-mass BBHs in the millihertz band  [^3] Note that for other kinds of bursting compact object binary, such as Double White Dwarfs (DWDs), there exists considerable uncertainty in their orbital evolution due to tidal interaction and mass transfer [see, e.g., 149, 150], particularly for cases involving extreme eccentricity. Therefore, our discussion is limited to BBHs, and we leave the discussion of other populations for future work. Nevertheless, it’s worth noting that the tidal interaction of DWDs tends to circularize the orbit, preventing eccentricity from reaching extreme values [see, e.g., 151, 152]. Additionally, the close approach of bursting DWDs at their pericenter may induce mass transfer and spin interaction, further restricting the lifetime of highly eccentric sources [see, e.g., Fig.4 in 33]. Thus, we suspect a limited population of highly eccentric DWDs and a negligible noise background from bursting DWDs in the millihertz band, even in the presence of external dynamical perturbations.. We consider three regimes that are expected to host eccentric BBHs: Globular Clusters (GCs), Galactic Field, and Galactic Nuclei (GNs). Within each of these regimes, we split the discussion into several specific cases specified below.
@@
In the simulation of BBHs in Globular Clusters and the Galactic Field, we adopt the steady-state approximation, assuming a continuous birth and death of compact object binaries in the universe while keeping the total number of GW sources unchanged. We calibrate our GW power spectrum based on the BBHs merger rate in the universe: $\Gamma_{\rm GC}=7.2_{-5.5}^{+21.5}\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$ [e.g., 108, 45, 109], $\Gamma_{\rm Field,Fly-by}=5_{-2}^{+5}\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$ [51]. Moreover, adopting the galaxy number density of $0.02~{}\rm Mpc^{-3}$ [110], we can use the BBH merger rate in the universe to estimate the merger rate in an averaged galaxy: $\gamma_{\rm GC}=3.6_{-2.7}^{+10.8}\times 10^{-7}\,\mathrm{yr}^{-1}$, $\gamma_{\rm Field,Fly-by}=2.5_{-1}^{+2.5}\times 10^{-7}\,\mathrm{yr}^{-1}$.
@@
For the Galactic Nuclei, we consider not only the old population of stars (calibrated using the steady-state approximation and the $m-\sigma$ relation [111, 112]), but also incorporate the non-equilibrium star formation history in the case of the Milky Way center, based on observational results [see, e.g., 113, 114, 115, 116].
@@
We note that the merger rate of BBHs can depend on the star formation history and exceed the estimation in the local universe at high redshift. This phenomenon is especially significant for the globular cluster channel [see, e.g., 117, 118, 119, 120, 121, 122]. Therefore, when calculating the cosmological background of bursting GW sources from GCs, we additionally take into account the change of BBH merger rate as a function of redshift and integrate Equation (21) up to redshift $z=3$ (i.e., $11.5$Gyr from present), following the merger rate evolution in figure 1 of Ref. [121]. The adopted redshift cutoff is partly justified because $z=3$ represents the epoch of the peak of globular cluster formation rate [118].
@@ III.2.1 Dynamically formed BBHs in Globular Clusters
Globular clusters are proposed to be a primary channel for the formation of GW mergers. In particular, the dense stellar environment within globular clusters results in various dynamical interactions, including scattering, few-body captures, the Eccentric Kozai-Lidov Mechanism, and Non-EKL triple interactions [see, e.g., 123, 40, 124, 44, 125, 43, 45]. Due to these frequent dynamical interactions, BBHs in globular clusters are anticipated to have non-negligible eccentricity, even within the frequency band of LIGO [see, e.g., 40, 44, 41, 42, 43, 45]. Moreover, since globular clusters widely exist in the universe (e.g., approximately 150 in the Milky Way [126, 127]), we expect a significant number of bursting sources originating from GCs.
For the GW background calculation, we adopt the BBHs eccentricity and semi-major axis distribution from previous studies [see, e.g., figure 4 in 43, for a summary], and assume that all the BBHs have the mass of $10-10~{}\rm M_{\odot}$ for simplicity. Using the steady-state approximation, we further average the power spectrum in Equation (20) over the evolution time of GW sources, i.e.,
$S_{n}(f)=\sum_{\rm sources}\int\left[\frac{2}{n(f)}\sqrt{g(n(f),e)}h_{0}(a)\right]^{2}\frac{1}{f_{\rm orb}}\frac{dt}{\tau_{\rm merger}}\ ,$ (22)
in which $\tau_{\rm merger}$ is the merger timescale of a BBH system.
We note that, in Equation (22), $e$, $a$, and $f_{\rm orb}$ are functions of time $t$. As a result, each integral gives the averaged GW power spectrum contribution of a bursting source throughout its lifetime. In the simulation, we initiate each system with the parameters at its formation, evolve the system until it merges, and adjust the collective GW power spectrum by varying the number of systems generated in the MCMC simulation. This adjustment ensures that the product of the number of systems and the inverse of their average lifetime equals the overall merger rate (see Section III.1).
@@
Figure 2: Total GW background from bursting BBHs in Globular Clusters and Galactic Field. Here we show (in dashed lines) the maximum collective GW power spectrum, $\sqrt{fS_{n}(f)}$, corresponding to bursting BBHs in the GCs (Left Panel) and Galactic Field (Right Panel) of Milky Way, M31, M33, and cosmological sources. The blue-colored region shows the total GW background from all four channels mentioned above, with its boundary representing the upper and lower limits. For comparison purposes, we also plot the LISA noise curve (N2A5 configuration, in red; eLISA, in blue [88, 89])
Figure 2, Left Panel, shows the maximum stochastic background from bursting BBHs in the Milky Way GCs (red dashed line), M31 GCs (green dotted line), M33 GCs (brown dashed-dot line), and GCs at a cosmological distance (blue dotted line). The overall GW power spectrum (depicted by the blue-shaded band) is contracted by adding all the sources, for which the maximum (minimum) value corresponds to the maximum (minimum) aforementioned merger rate. As can be seen in Left Panel of Figure 2, bursting BBHs within Milky Way GCs dominate the overall GW power spectrum. Meanwhile, the contributions from the other three channels have comparable levels, approximately two orders of magnitude weaker than Milky Way GCs. Moreover, the total GW background reaches the LISA noise curve at around 5 mHz, which suggests that neglecting these sources in future data analysis could significantly compromise LISA’s sensitivity.
We note that, previous studies [see, e.g., 74, 73, 40] have shown the overall cosmological GW background from stellar-mass BBHs, accounting for all the eccentricity values, is either comparable to or smaller than the millihertz LISA noise level. For our cases, as the highly eccentric binaries represent a subset of the entire BBH population, we anticipate their cosmological GW spectrum level to be lower than the overall background from BBHs and, thus fall below the LISA noise curve. This prediction serves as an upper bound for the cosmological background of bursting sources in the simulation, and turns out to be consistent with the simulation result (depicted by the blue dotted line in Figure 2).
@@ III.2.2 Fly-by induced, highly eccentric BBHs in the Galactic Field
Recent studies show that external perturbations, such as flybys or galactic tides, can exert a significant influence on the orbital evolution of wide compact object binaries in the galactic field [see, e.g., 129, 51, 52, 53]. Although the isolated wide binaries are not likely to merge due to their large GW merger timescale (sometimes exceeding the Hubble time, as indicated by Equation (9)), the dynamical perturbations from the environment can drive their eccentricity to extreme values, making the BBH system emit GW bursts during pericenter passage, and resulting in a GW merger. Given the substantial number of stars in the galactic field, the fly-by-induced bursting BBHs can significantly contribute to the overall GW background, even if their fraction within the entire stellar population is small.
We follow a similar approach for the GW background calculation as shown in Section III.2.1. In particular, we generate the spacial distribution of wide BBHs according to the density profile of a Milky-Way type galaxy (see Eq. (23$\sim$(26) in Ref. [51]). We randomly generate the semi-major axis of BBHs with the log-uniform distribution (from $100$ to $50000~{}\rm au$), then calculate the fly-by merger rate as a function of semi-major axis [see Eq. (16$\sim$(23) in 53]. With a given semi-major axis, there is a critical eccentricity, $e_{\rm crit}$, for a BBH system to merge before the next fly-by [see Eq. (2) in 53]. Therefore, we randomly generate the initial eccentricity of fly-by induced merger in the range of $e_{\rm crit}$ to $1$, following the thermal distribution $F(e)=2e$.
@@
III.2.3 BBHs in Galactic Nuclei
The galactic nucleus, particularly if hosting an SMBH, can be a promising environment for the formation of stellar-mass bursting BBHs [see, e.g., 38, 108, 58, 130, 131, 60, 63, 132, 133]. For example, consider a stellar-mass BBH system orbiting the SMBH in the galactic nucleus. The EKL mechanism can induce eccentricity excitation in the inner binary’s orbit, potentially driving the eccentricity to extreme values and triggering GW bursts [57, 56, 58, 59, 60, 61, 62, 63]. Consequently, we anticipate a high fraction of bursting BBHs in the galactic center, especially within the Nuclear Star Cluster, where the influence of the SMBH dominates the dynamical environment.
@@ III.2.3 BBHs in Galactic Nuclei
However, observation shows there is a young nuclear star cluster (YNC) at the Milky Way center, which is formed approximately $2\sim 8~{}\rm\,Myr$ ago [see, e.g., 113, 134, 115, 116]. Moreover, there might be a hidden mass of $\sim 3000~{}\rm M_{\odot}$ within the inner $1000$ au of the Milky Way center [e.g., 135, 136], supported by the observation of stellar motion and the theoretical arguments of the stability of S0-2, [e.g., 137, 138, 139, 140, 141]. In our previous work [33], we carried out Monte-Carlo simulations of these BBHs’ evolution, taking into account the EKL effect up to the octupole level of approximation [142], general relativity precession [e.g., 143], and GW emission [17]. The results indicated that the young population of stars potentially has a much higher fraction of bursting BBHs, especially when the YNC’s age is below the detectable lifetime of bursting sources [see, e.g., Fig.7 in 33].

### v2

@@ II.3.3 Average Stochastic Background from Extragalactic Sources
here $\dot{\varepsilon}(f_{\mathrm{em}},z)\Delta f_{\mathrm{em}}$ represents the rate (per unit proper time, per unit co-moving volume) at which GW energy between frequencies $f_{\mathrm{em}}$ and $f_{\mathrm{em}}+\Delta f_{\mathrm{em}}$ is emitted into the universe, at red-shift z. $f_{\mathrm{em}}=f(1+z)$ denotes the emitted GW frequency (i.e., the frequency measured by a contemporaneous observer, not the redshifted frequency $f$ we measure today). We note that Equation (25) has taken into account the decrease in the waves’ energy due to the redshift. For a detailed explanation, see Eqs. 34 - 35 in Barack and Cutler [74].
Furthermore, we calculate $\dot{\varepsilon}(f_{\mathrm{em}},z)$ using the following relation:
$\dot{\varepsilon}(f_{\mathrm{em}},z)=\Gamma(z)\left\langle\frac{dE_{\mathrm{GW}}\left(f_{\mathrm{em}}\right)}{df_{\mathrm{em}}}\right\rangle.$ (26)
In Equation (26), $\Gamma(z)=dN/(dtdV)$ represents the comoving merger rate of bursting BBHs per unit proper time, per unit co-moving volume; $\left\langle dE_{\mathrm{GW}}\left(f_{\mathrm{em}}\right)/df_{\mathrm{em}}\right\rangle$ is the emitted GW energy per unit frequency from a bursting BBH system, averaged over different BBH systems’ initial conditions from the population synthesis.
We note that, the meaning of $dE_{\rm GW}/df_{\mathrm{em}}$ in Equation (26) is different from $d\rho_{\rm single}/df_{\mathrm{em}}$ in Equation (20). In particular, $E_{\rm GW}$ accounts for the total gravitational wave energy emitted throughout the evolution of a source, while $\rho_{\rm single}$ represents the source’s gravitational wave energy density at a given location in space (or equivalently, the energy flux $F_{\rm single}=c\rho_{\rm single}$, where $c=1$ is the speed of light). These two quantities can be related by integrating the energy flux $F$ on a spherical surface $\Omega_{s}$, with a fixed radius of $d_{l}$ from the BBH system, throughout the sources’ time evolution. (see, e.g., Eq. 25 in Phinney [110])
[…]
and thus:
[…]
In Equation (28), $F_{\rm single}=c\rho_{\rm single}$ can be expressed as a function of $f_{\rm em}$ and $t$. This is because, at any given time during the system’s evolution, $\rho_{\rm single}$ is a function of the emitted GW frequency (see Equation (20)). However, over the long-term evolution, other orbital parameters also become functions of time (see the explanation below Equation (22)). Therefore, when integrating over the time evolution of a GW source, Equation (28) yields the cumulative energy emitted within a fixed frequency bin.
We note that, Eq. 25 in Phinney [110] includes the effect of redshift at arbitrarily large luminosity distance, while here we drop the term of $1+z$ in Equation (27). This is because our purpose is to get the total energy measured as a function of the emitted frequency in the comoving frame of the source (the cosmological effects have been taken into account via Equation (25)). In other words, we choose the spherical surface $\Omega_{s}$ with a short distance $d_{l}$ from the BBH system, such that the effect of redshift is negligible when we sum over the energy flux. In fact, the factor $d_{l}^{2}$ will cancel out with the amplitude term in $\rho_{\rm single}$ on the right side of Equation (28), thus its value will not affect the calculation of $dE_{\rm GW}/df_{\rm em}$, since the energy density $\rho_{\rm single}$ is proportional to the square of the strain amplitude (see Equation (20)), while the strain amplitude is inversely proportional to the distance (see Equations (13) - (14)).
We emphasize that, in general, there exists a straightforward relationship between the spectrum of the gravitational wave background produced by a cosmological distribution of discrete gravitational wave sources and the present-day comoving number density of remnants, as described by Eq. 5 in Ref. [110]. Realistic examples and further discussions on this relationship can be found in studies such as Refs. [111, 108]. For our cases, however, since the local population can have a non-negligible contribution to the overall GW background level, we include the inhomogeneous spatial distribution and non-equilibrium star formation history in the Monte Carlo simulations detailed below (see Section III.1). Therefore, for consistency, we adopt Equations (21), (24) and (25) as a robust approach to estimate the GW background level across all channels, including cosmological background calculations.
@@ III.1 Configuration of the Simulations
As a proof of concept, we adopt the population model in Xuan et al. [33] to generate the parameters of highly eccentric, stellar-mass BBHs in the millihertz band  [^3] Note that for other kinds of bursting compact object binary, such as Double White Dwarfs (DWDs), there exists considerable uncertainty in their orbital evolution due to tidal interaction and mass transfer [see, e.g., 155, 156], particularly for cases involving extreme eccentricity. Therefore, our discussion is limited to BBHs, and we leave the discussion of other populations for future work. Nevertheless, it’s worth noting that the tidal interaction of DWDs tends to circularize the orbit, preventing eccentricity from reaching extreme values [see, e.g., 157, 158]. Additionally, the close approach of bursting DWDs at their pericenter may induce mass transfer and spin interaction, further restricting the lifetime of highly eccentric sources [see, e.g., Fig.4 in 33]. Thus, we suspect a limited population of highly eccentric DWDs and a negligible noise background from bursting DWDs in the millihertz band, even in the presence of external dynamical perturbations.. We consider three regimes that are expected to host eccentric BBHs: Globular Clusters (GCs), Galactic Field, and Galactic Nuclei (GNs). Within each of these regimes, we split the discussion into several specific cases specified below.
@@
In the simulation of BBHs in Globular Clusters and the Galactic Field, we adopt the steady-state approximation, assuming a continuous birth and death of compact object binaries in the universe while keeping the total number of GW sources unchanged. We calibrate our GW power spectrum based on the BBHs merger rate in the universe: $\Gamma_{\rm GC}=7.2_{-5.5}^{+21.5}\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$ [e.g., 113, 45, 114], $\Gamma_{\rm Field,Fly-by}=5_{-2}^{+5}\,\mathrm{Gpc}^{-3}\,\mathrm{yr}^{-1}$ [52]. Moreover, adopting the galaxy number density of $0.02~{}\rm Mpc^{-3}$ [115], we can use the BBH merger rate in the universe to estimate the merger rate in an averaged galaxy: $\gamma_{\rm GC}=3.6_{-2.7}^{+10.8}\times 10^{-7}\,\mathrm{yr}^{-1}$, $\gamma_{\rm Field,Fly-by}=2.5_{-1}^{+2.5}\times 10^{-7}\,\mathrm{yr}^{-1}$.
@@
For the Galactic Nuclei, we consider not only the old population of stars (calibrated using the steady-state approximation and the $m-\sigma$ relation [116, 117]), but also incorporate the non-equilibrium star formation history in the case of the Milky Way center, based on observational results [see, e.g., 118, 119, 120, 121].
@@
We note that the merger rate of BBHs can depend on the star formation history and exceed the estimation in the local universe at high redshift. This phenomenon is especially significant for the globular cluster channel [see, e.g., 122, 123, 124, 125, 126, 127]. Therefore, when calculating the cosmological background of bursting GW sources from GCs, we additionally take into account the change of BBH merger rate as a function of redshift and integrate Equation (25) up to redshift $z=3$ (i.e., $11.5$Gyr from present), following the merger rate evolution in figure 1 of Ref. [126]. The adopted redshift cutoff is partly justified because $z=3$ represents the epoch of the peak of globular cluster formation rate [123].
@@ III.2.1 Dynamically Formed BBHs in Globular Clusters
Globular clusters are proposed to be a primary channel for the formation of GW mergers. In particular, the dense stellar environment within globular clusters results in various dynamical interactions, including scattering, few-body captures, the Eccentric Kozai-Lidov Mechanism, and non-secular triple interactions [see, e.g., 128, 40, 129, 44, 130, 43, 45]. Due to these frequent dynamical interactions, BBHs in globular clusters are anticipated to have non-negligible eccentricity, even within the frequency band of LIGO [see, e.g., 40, 44, 41, 42, 43, 45]. Moreover, since globular clusters widely exist in the universe (e.g., approximately 150 in the Milky Way [131, 132]), we expect a significant number of bursting sources originating from GCs.
For the GW background calculation, we adopt the BBH eccentricity and semi-major axis distribution from previous studies [see, e.g., figure 4 in 43, for a summary], and assume that all the BBHs have the mass of $10-10~{}\rm M_{\odot}$ for simplicity. Since we are interested in the bursting stage where BBHs are in a wide configuration, we evolve the systems shown in Martinez et al. [43], which is in a higher frequency band, back to the former bursting stage. With the aforementioned parameter distribution, we carry out Monte Carlo simulations, randomly generate BBH systems, and evolve them until the merger happens. The stochastic background is calculated under the steady-state approximation, using Equation (24).
@@
Figure 2: Total GW background from bursting BBHs in Globular Clusters and Galactic Field. Here we show (in dashed lines) the maximum collective GW power spectrum, $\sqrt{fS_{n}(f)}$, corresponding to bursting BBHs in the GCs (Left Panel) and Galactic Field (Right Panel) of Milky Way, M31, M33, and cosmological sources. The blue-colored region shows the total GW background from all four channels mentioned above, with its boundary representing the upper and lower limits. For comparison purposes, we also plot the LISA noise curve (N2A5 configuration [92, 93, 134], in the solid red line).
Figure 2, Left Panel, shows the maximum stochastic background from bursting BBHs in the Milky Way GCs (red dashed line), M31 GCs (green dotted line), M33 GCs (brown dashed-dot line), and GCs at a cosmological distance (blue dotted line). The overall GW power spectrum (depicted by the blue-shaded band) is obtained by summing the contributions of all the sources mentioned above, where the maximum (minimum) value corresponds to the maximum (minimum) aforementioned merger rate. As can be seen in Left Panel of Figure 2, bursting BBHs within Milky Way GCs dominate the overall GW power spectrum. Meanwhile, the contributions from the other three channels have comparable levels, approximately two orders of magnitude weaker than Milky Way GCs. Moreover, the total GW background reaches the LISA noise curve at around 5 mHz, which suggests that neglecting these sources in future data analysis could significantly compromise LISA’s sensitivity.
We note that previous studies [see, e.g., 77, 76, 40] have shown the overall cosmological GW background from stellar-mass BBHs, accounting for all the eccentricity values, is either comparable to or smaller than the millihertz LISA noise level. For our cases, as the highly eccentric binaries represent a subset of the entire BBH population, we anticipate their cosmological GW spectrum level to be lower than the overall background from BBHs and, thus, fall below the LISA noise curve. This prediction serves as an upper bound for the cosmological background of bursting sources in the simulation and turns out to be consistent with the simulation result (depicted by the blue dotted line in Figure 2).
@@ III.2.2 Fly-by Induced, Highly Eccentric BBHs in the Galactic Field
Recent studies show that external perturbations, such as flybys or galactic tides, can exert a significant influence on the orbital evolution of wide compact object binaries in the galactic field [see, e.g., 135, 52, 53, 54]. Although the isolated wide binaries are not likely to merge due to their large GW merger timescale (sometimes exceeding the Hubble time, as indicated by Equation (9)), the dynamical perturbations from the environment can drive their eccentricity to extreme values, making the BBH system emit GW bursts during pericenter passage, and resulting in a GW merger. Given the substantial number of stars in the galactic field, the fly-by-induced bursting BBHs can significantly contribute to the overall GW background, even if their fraction within the entire stellar population is small.
We follow a similar approach for the GW background calculation as shown in Section III.2.1. In particular, we generate the spacial distribution of wide BBHs according to the density profile of a Milky-Way type galaxy (see Eqs. 23 - 26 in Ref. [52]). We randomly generate the semi-major axis of BBHs with the log-uniform distribution (from $100$ to $50000~{}\rm au$), then calculate the fly-by merger rate as a function of semi-major axis [see Eqs. 16 - 23 in 54]. With a given semi-major axis, there is a critical eccentricity, $e_{\rm crit}$, for a BBH system to merge before the next fly-by [see Eq. 2 in 54]. Therefore, we randomly generate the initial eccentricity of fly-by induced merger in the range of $e_{\rm crit}$ to $1$, following the thermal distribution $F(e)=2e$.
@@
III.2.3 BBHs in the Galactic Nuclei
The galactic nucleus, particularly if hosting an SMBH, can be a promising environment for the formation of stellar-mass bursting BBHs [see, e.g., 38, 113, 59, 136, 137, 61, 64, 138, 139]. For example, consider a stellar-mass BBH system orbiting the SMBH in the galactic nucleus. The EKL mechanism can induce eccentricity excitation in the inner binary’s orbit, potentially driving the eccentricity to extreme values and triggering GW bursts [58, 57, 59, 60, 61, 62, 63, 64]. Consequently, we anticipate a high fraction of bursting BBHs in the galactic center, especially within the Nuclear Star Cluster, where the influence of the SMBH dominates the dynamical environment.
@@ III.2.3 BBHs in the Galactic Nuclei
However, observation shows there is a young nuclear star cluster (YNC) at the Milky Way center, which is formed approximately $2\sim 8~{}\rm\,Myr$ ago [see, e.g., 118, 140, 120, 121]. Moreover, there might be a hidden mass of $\sim 3000~{}\rm M_{\odot}$ within the inner $1000$ au of the Milky Way center [e.g., 141, 142], supported by the observation of stellar motion and the theoretical arguments of the stability of S0-2, [e.g., 143, 144, 145, 146, 147]. In our previous work [33], we carried out Monte-Carlo simulations of these BBHs’ evolution, taking into account the EKL effect up to the octupole level of approximation [148], general relativity precession [e.g., 149], and GW emission [17]. The results indicated that the young population of stars potentially has a much higher fraction of bursting BBHs, especially when the YNC’s age is below the detectable lifetime of bursting sources [see, e.g., Fig.7 in 33].